OpenStudy (anonymous):

using the formula C = −3000 ln(x/y) where y is the intensity of the incident light and x is the intensity of light that emerges. Find the concentration of the substance if the intensity x is 30% of y.

zepdrix (zepdrix):

Find C when `x is 30% of y`. When \(\large x=.30y\) Let's solve for \(\large \dfrac{x}{y}\). Dividing both sides by y gives us,\[\large \color{orangered}{\frac{x}{y}=.30}\] Substitute this into the equation and solve for C! :) \[\large C=-3000\ln\left(\color{orangered}{\frac{x}{y}}\right)\]

OpenStudy (anonymous):

wow, that's colorful